# The Artificial Scientist: in-transit Machine Learning of Plasma Simulations

### v2

@@ IV-C Machine learning
The task of in-transit ML is to extract patterns from the live observations and thereby learn correlations between variables and reduced representations. Here, we focus on predicting particle dynamics from radiation spectra, which constitutes an ill-posed problem in that any given radiation spectrum can potentially be produced by various types of dynamics. When learning in-transit from a simulation, this problem is reduced from an intractable general case to the question of: which type of local particle dynamics under the given physical setup, i. e. the relativistic KHI, gave rise to an observed radiation spectrum observed in one direction, which, however, remains an ill-posed problem. Many ML techniques, especially feed-forward networks, tend to predict the mean of possible inversions, even if this itself is not a valid solution [43]. One class of techniques specifically designed to learn by sampling from conditional multi-modal distributions, such as those that occur as solutions to our ill-posed problem, are flow-based models [44, 45] and invertible neural networks (INNs) [46]. These learn an invertible transformation, or change-of-variables, between a chosen posterior distribution, often a normal distribution , and a prior described by the training data, in our case the distribution of observed particle dynamics . One specific difference between them is, that INNs also learn the deterministic forward function defining the inverse problem, in our case the mapping between samples of particle dynamics and radiation spectrum . Put differently, they also learn to predict the condition under which a given sample of the prior may occur.
@@
Particle configurations in phase space of the plasma dynamics simulated here can be very complex, such that a small sample of particles cannot adequately represent the dynamics. For INNs and flow-based models this poses a challenge, because the change-of-variables approach [47, 48] dictates, that the information volume, i. e. the number of activations at the input and output of each coupling block, must be constant throughout the network, which results in a very large network when working with a prior distribution of dimensionality on the order of . In this setting, the number of network parameters can be kept in check by using affine flows [49, 44] with designed convolutional sub-networks which should also be invariant to transposition of particles in the input vector to be efficiently trainable.
@@
When generating point cloud (PC), flow-based models are typically tasked with sampling them point-by-point, i.e. the model learns a probability density that describes the particle density in phase space. The main drawback here is, that the model has to learn the many-modal distributions which can occur in the KHI phase space on a single-particle basis and can not longer generate multiple particle distribution for a single condition, i.e. not distinctly produce all PCs solving an ill-posed problem. Another issue is, that, unless we use an encoder (conditioning network) producing very small latent vectors of around six dimensions for the radiation, the dimension of the condition is much larger than the dimension of the prior while the network has to be sampled times to retrieve a useful particle distribution.
@@
To overcome the above inefficiencies, we opted to view the particle configuration as a point cloud (PC) and reduce the dimensionality of the prior by tasking an outer network to learn latent representations of the phase space, which are also invariant with respect to transposition of particles. This outer network can in principle be an autoencoder (AE) whose latent representation is passed to the INN as the prior . In practice, the INN will not generate exact latent vectors on its backward pass, but variations, which will cause the AE’s decoder to fail at reconstruction. Therefore, we employ a variational auto encoder (VAE) instead, which, albeit harder to train, intrinsically trains its decoder to be robust against the variations produced by the INN.
@@
The encoder (light green) follows an architecture proposed in [50], which by itself is a simplified version of PointNet [51], using convolutions acting on the coordinates of particles separately to gradually (channels: ) extract 608 features per particle, followed by a max-pooling to obtain a transposition-invariant feature set, which is then processed by two multi-layer perceptrons (MLP), each with hidden features, to obtain predictions for the mean and standard deviation to sample the 544-dimensional latent vector .
@@
As the reconstruction loss of the VAE, we use Chamfer’s distance (CD) [53], which is the most commonly used differentiable function to compare PCs in literature, because it is efficient to compute.
Regularization of the VAE’s latent space is controlled by Kullback-Leibler divergence (KL) [54].
The resulting spectrum of the INNs forward function is compared to the ground truth using mean-squared error
The posterior normal distribution, produced by the INN’s forward pass, and the prior distribution, produced by the INN’s backward pass, are both evaluated using maximum mean discrepancy (MMD) using an inverse multi-quadratic kernel, as proposed in [46].
The total loss corresponds to the weighted sum
@@
While training on the continuous data stream of non-steady configurations, we employ experience replay (EP) [55] to avoid catastrophic forgetting of earlier simulation time steps while training on later ones. To integrate with the streaming setup, we implemented this approach to continual learning as a separate component, called training buffer, between the streaming receiver and the training loop. This component holds two buffers: a now-buffer, which holds the latest samples received. Each time new samples are received, they are prepended to the now-buffer and any element which would then exceed the buffer size is removed and added to the EP buffer. The EP buffer holds a maximum of elements. When new samples are to be added to the EP buffer when it is full, a randomly chosen item is removed. For a single iteration of the training loop, a batch of data is requested from the training buffer. This batch constitutes the training batch. It is generated by taking random samples form the now-buffer and random samples from the EP buffer, yielding a training batch size of . In order to increase training success, we perform iterations of the training loop per single time step from the data stream. For each iteration, a new training batch is generated from the now- and EP buffer.
@@
For all training runs, we use the Adam optimizer [56] with , , and an weight decay strength of . For data parallel training we employ PyTorch’s Distributed Data Parallel (DDP) module with its N/RCCL backend. Learning rates are scaled following a square-root rule [57]. In order to accelerate learning with large batch sizes, we train layers belonging to the VAE at a learning rate higher by a factor than that of the INN.
Technical Challenges: The main drawback of CD, used as the VAE’s reconstruction loss, is its lacking sensitivity to point density. It also fails to distinguish some point configurations [58]. Both of these problems are overcome by the earth mover’s distance (EMD) [59] loss, but at much higher computational cost111We observed about a increase in batch times when using EMD [60] vs. a simple implementation of CD.. Still, a rather compute and memory efficient differentiable implementation of EMD is available in the geomloss python package [60], which, instead of relying entirely on PyTorch, internally uses the KeOps library [61, 62] for performance. Unfortunately, the KeOps library is implemented directly in CUDA, and thus cannot be used on AMD GPUs in Frontier. Hence, we were barred from employing EMD loss in our runs on Frontier. Perhaps the community needs a HIP version of the KeOps library.
@@ IV-D Setup on ORNL Frontier system
While distributing nodes exclusively to either PIConGPU or to the MLapp, inter-node setup in Fig. 3(c), is easier to achieve in Slurm (and this strategy would be supported by our flexible I/O solution), we decide to use the intra-node setup in Fig. 3(c) and share each compute node between PIConGPU and the MLapp. This has the advantage that data exchange mostly needs not leave the node, but requires a more specific resource allocation: Each single compute node no longer runs homogeneous parallel sub-jobs, but has to locally assign resources to heterogeneous applications. For our setup, this local assignment gives 4 GCDs to PIConGPU and 4 GCDs to MLapp. In openPMD and ADIOS2, each reader application decides on its own which remote datasets to load. Here, the distribution of loaded data regions (i. e. local blocks within the global dataset) is configured such that data is shared within node boundaries.
@@
Due to the plugin-based structure of PIConGPU, both the particle and radiation input required by MLapp are provided by distinct output plugins of PIConGPU, meaning that two parallel data streams are opened between PIConGPU and the MLapp. Each plugin prepares the current time step’s data and, upon having all data ready, ADIOS2 will gather the metadata, including remote read addresses, to parallel rank 0 and provide the time step to the reader. The reader performs remote reads to its liking and closes the time step, indicating to the writer that the data can now be dropped.
@@
A unique and important requirement of the presented workflow is the ability to not only run but also scale on a large number of Frontier’s AMD MI250X GPUs. Our machine learning model is small enough to fit on a single GCD; therefore, parallel training of this model is done using data parallelism where copies of the model are distributed across GCDs with each copy of the model receiving different chunks of data to train on. Once each model computes its gradients, all the instances of the model must do a collective all-reduce communication to average the gradients. We expect the scaling of the workflow to critically depend on the optimization of this all-to-all communication in PyTorch DDP, since communication within the PIConGPU simulation is only between next neighbors. We experienced a hard limit to scaling in our current setup: The all-to-all communication between PyTorch DDP ranks using the N/RCCL backend hits system limitations on the possible number of open sockets beyond 100 nodes. Potential options to circumvent this limit include using a libfabric backend for N/RCCL or using PyTorch DDP’s MPI backend. While the N/RCCL backend was able to support PyTorch up to the scale used in this study, we can already conclude that further extensive evaluation of the communication backends within PyTorch will become mandatory to provide a stable base for our near future scaling studies on Frontier. This area of research appears to be still under development, to the best of our knowledge [63].
@@ V-A Weak-Scaling
When training any kind of statistical model, in our case a deep neural network, one has to ensure that for each characteristic of the domain one hopes to learn, there are sufficient examples available in the dataset. When the dataset is provided in-transit, redundancy must still be built into the setup, either by using a large simulation volume or running multiple independent simulations concurrently, to allow the network to learn repeated patterns of the simulations and thereby distinguish relevant domain features, such as physical relationships, from noise. For this reason, in this work, we focus on a weak-scaling study of the proposed pipeline, as we anticipate this to be the most likely use-case. To evaluate parallel scaling as well as learning scaling, we performed simulations with in-transit training from 8 up to 96 nodes. The lower limit of 8 nodes is set by physical size constraints of the simulation. Fig. 8 shows the weak scaling behavior of the pipeline on Frontier. We define efficiency as the ratio of runtime at a specific size over runtime of the smallest size. Before averaging the times per training batch, we removed outliers more than four standard-deviations from the mean, because on Frontier we observed single batches taking more than the mean time.
While ideal scaling should correspond to 100 % efficiency across all sizes, communication overheads reduce efficiency in practice. Here, the efficiency of in-transit training reaches around 35 % at 96 nodes. We attribute the low parallel efficiency to two main reasons: One of them is the inevitable all-to-all communication between PyTorch ranks that takes place to average the gradients during each backward pass and obtain a synchronized result on all ranks, which accounts for a deficit of , which is still a respectable result achieved by overlapping computations and communication in PyTorch DDP. The other is the computation of the two MMD loss terms, which amount to matrix dot products with data distributed across all ranks. Due to torch not offering a distributed primitive and the matrices involved being relatively small, even at the largest achieved scale, we opted for a naive implementation which replicates work between nodes. The global communication required here222torch.distributed.all_gather_into_tensor breaks the torch computational graph, i.e. synchronizes graph execution with host code at the invocation site. To our knowledge there is no high-level API available to stitch these operations into the graph to allow them to be executed asynchronously to both the host and each other, i.e no enqueueing of host functions or events.
@@ V-A1 Scaling Model Training
In our scaling runs, we explored training from scratch at large scale, sampling up to batches per sample consumed from the stream and found learning success up to about . We settled on a base-learning rate of , scaled according to the square-root scaling rule [57] depending on the actual number of GCDs. We observe a training speed up with larger batch sizes. That is, the increased learning rate overcompensates the training slow down due to larger batch sizes. Furthermore, it became evident that the three losses connected to the INN converged best at lower learning rates, while the VAE only found better minima at the highest learning rate, showing that separate learning rates and of the encoder/decoder block and the inversion block, respectively, need to be applied at large scales.
Within the scope of this study, we have to conclude, that for in-transit training at very large batch sizes, hyper parameter studies do not transfer from small-scale experiments and thus have to be performed at scale. To enable in-transit learning at the scales we tested or beyond, comprehensive studies of the relations between the block learning rates and , batch sizes, and maybe even loss weights in Eq. (1) have to be performed. Exploring this parameter space for the large scale training requires the full study to be systematically performed at these scales (up to 384 GCDs), either by employing a streamed simulation or the same training data stored on disk. We argue, that the latter is challenging at scale due to the lack of disk capacity and bandwidth.
@@ V-B Quantifying the predictive capabilities of the machine learning method
The momentum prediction from the KHI shear surface was expected to be more difficult since particles from both directions cross in the shear surface, resulting in two separate and distinct particle populations. The ML reconstruction consistently predicts these two distinct populations. This is consistent with the interpretation, that the INN has learned unsupervised that the vortex region undergoes more acceleration and thus emits more intense radiation than the bulk of the plasma. However, the momentum prediction is less accurate: the mean of (one of) these populations often does not match the PIConGPU data (Fig. 9(b vs. c) green). We suspect that this is partly due to our choice of region boundaries on the shear surface, which results in a much lower sampling rate of one flow direction compared to the other, leading to a lower accuracy in predicting the mean of the lower peak of the momentum distribution. It is also possible, that this problem arises from the VAE’s decoder having difficulties reconstructing discontinuous distributions. Nevertheless, the successful prediction of shear regions indicates that the encoder extracts crucial features from the input data to recognize whether a region contains a vortex and encodes them in latent space.
Despite suboptimal reconstruction of vortex momenta, the ML prediction still allows to clearly identify KHI regions. Previous work on PC generation [64] observed that the latent space of an AE may contain more information than the decoder can reconstruct with satisfactory fidelity and that more sophisticated generative models, such as GANs or flows, can learn to produce more detailed reconstructions when conditioned on the same latent space.

### v3

@@ IV-C Machine learning
The task of in-transit ML is to extract patterns from the live observations and thereby learn correlations between variables and reduced representations. Here, we focus on predicting particle dynamics from radiation spectra, which constitutes an ill-posed problem – since any given radiation spectrum can potentially be produced by various types of dynamics [28]. When learning in-transit from a simulation, this problem is reduced from an intractable general case to the question of: which type of local particle dynamics under the given physical setup, i. e. the relativistic KHI, gave rise to an observed radiation spectrum observed in one direction, which, however, remains an ill-posed problem [28]. Many ML techniques, especially feed-forward networks, tend to predict the mean of possible inversions, even if this itself is not a valid solution [46]. One class of techniques specifically designed to learn by sampling from conditional multi-modal distributions, such as those that occur as solutions to our ill-posed problem, are flow-based models [47, 48] and invertible neural networks (INNs) [49]. These learn an invertible transformation from a chosen posterior distribution (often a normal distribution ) and a prior described by the training data (in our case the distribution of observed particle dynamics ) and vice versa. One specific difference between them is that INNs also learn the deterministic forward function defining the inverse problem – in our case, the mapping between samples of particle dynamics and radiation spectrum . Put differently, INNs also learn to predict the condition for a given prior .
@@
Particle configurations in phase space of the plasma dynamics simulated here can be very complex, such that a small sample of particles cannot adequately represent the dynamics. For INNs and flow-based models this poses a challenge, because the change-of-variables approach [50, 51] dictates that the information volume, i. e. the number of activations at the input and output of each coupling block, must be constant throughout the network, which results in a very large network when working with a prior distribution of dimensionality on the order of . In this setting, the number of network parameters can be kept in check by using affine flows [52, 47] with convolutional sub-networks designed to be invariant to transposition of particles in the input vector to be efficiently trainable.
@@
When generating point cloud (PC), flow-based models are typically tasked with sampling them point-by-point, i.e. the model learns a probability density describing the particle density in phase space. The main drawback of this method is that the model has to learn the many-modal distributions which can occur in the KHI phase space on a single-particle basis and can no longer generate multiple particle distribution for a single condition, i. e. we cannot expect it to distinctly produce all PCs solving an ill-posed problem [28]. Another issue is that, unless we use an encoder (conditioning network) producing very small latent vectors of around six dimensions for the radiation, the dimension of the condition is much larger than the dimension of the prior while the network has to be sampled times to retrieve a useful particle distribution.
@@
To overcome the above inefficiencies, we opted to 1. view the particle configuration as a point cloud (PC), and 2. reduce the dimensionality of the prior by tasking an outer network to learn latent representations of the phase space that are also invariant with respect to transposition of particles. This outer network can, in principle, be an autoencoder (AE) whose latent representation is passed to the INN as samples of the prior distribution . In practice, the INN will not generate exact latent vectors on its backward pass, but variations, which will cause the AE’s decoder to fail at reconstruction. Therefore, we employ a variational auto encoder (VAE) instead, which, albeit harder to train, intrinsically trains its decoder to be robust against the variations produced by the INN.
@@
The encoder (light green) follows an architecture proposed in [53], which is a simplified version of PointNet [54]. 6-dimensional vectors corresponding to the positions and momenta of the particles are fed to convolutions separately to extract 608 features per particle (channels: ), followed by a max-pooling to obtain a transposition-invariant feature set. These 608 features are then processed by two multi-layer perceptrons (MLP), each with hidden features, to obtain predictions for the mean and standard deviation to sample the 544-dimensional latent vector .
@@
: As the reconstruction loss of the VAE, we use the Chamfer’s distance (CD) [56], it being the most popular differentiable function for comparing PCs the in literature, owing to it being cheap to compute.
: Regularization of the VAE’s latent space is controlled by Kullback-Leibler divergence (KL) [57].
: The resulting spectrum of the INNs forward function is compared to the ground truth using mean-squared error
, : Both the posterior normal distribution produced by the INN’s forward pass and the prior distribution produced by the INN’s backward pass are evaluated using maximum mean discrepancy (MMD) using an inverse multi-quadratic kernel [49].
The total loss corresponds to the weighted sum , where,
@@
While training on the continuous data stream of non-steady configurations, we employ experience replay (EP) [58] to avoid catastrophic forgetting of earlier simulation time steps while training on later ones. To integrate with the streaming setup, we implemented this approach to continual learning as a separate component, called training buffer, between the streaming receiver and the training loop. This component holds two buffers: a now-buffer, and an EP-buffer. The now-buffer holds the latest samples received. Each time new samples are received, they are prepended to the now-buffer and any element which would then exceed the buffer size is removed and added to the EP buffer. The EP buffer holds a maximum of elements. When new samples are to be added to the EP buffer when it is full, a randomly chosen item is removed. For a single iteration of the training loop, a batch of data is requested from the training buffer. This batch constitutes the training batch. It is generated by taking random samples form the now-buffer and random samples from the EP buffer, yielding a training batch size of . In order to increase training success, we perform iterations of the training loop per single time step from the data stream. For each iteration, a new training batch is generated from the now- and EP buffer.
@@
For all training runs, we use the Adam optimizer [59] with , , and weight decay . For data parallel training we employ PyTorch’s Distributed Data Parallel (DDP) module with its N/RCCL backend. Learning rates are scaled following a square-root rule [60]. In order to accelerate learning with large batch sizes, we train layers belonging to the VAE at a learning rate higher by a factor than that of the INN.
Technical Challenges: The main drawback of CD, used as the VAE’s reconstruction loss, is its lacking sensitivity to point density. It also fails to distinguish some point configurations [61]. Both of these problems are overcome by the earth mover’s distance (EMD) [62] loss, but at much higher computational cost111We observed about a increase in batch run times when using EMD [63] compared to a simple implementation of CD.. Still, a rather compute and memory efficient differentiable implementation of EMD is available in the geomloss python package [63], which, instead of relying entirely on PyTorch, internally uses the KeOps library [64, 65] for performance. Unfortunately, the KeOps library is implemented directly in CUDA, and thus cannot be used on AMD GPUs in Frontier. Hence, we were barred from employing EMD loss in our runs on Frontier. Perhaps the community needs a HIP version of the KeOps library.
@@ IV-D Setup on ORNL Frontier system
While distributing nodes exclusively to either PIConGPU or to the MLapp is easier to achieve in Slurm (i. e. inter-node setup in Fig. 3(c) that is supported by our flexible I/O solution), we decide to use the intra-node setup from Fig. 3(c) where each compute node is shared between PIConGPU and the MLapp. This strategy has the advantage that data exchange mostly does not need to leave the node, but requires a more specific resource allocation. More specifically, each single compute node no longer runs homogeneous parallel sub-jobs, but has to locally assign resources to heterogeneous applications. For our setup, this local assignment gives 4 Graphics Compute Dies (GCDs) to PIConGPU and 4 GCDs to MLapp. In openPMD and ADIOS2, each reader application decides on its own which remote datasets to load. Here, the distribution of loaded data regions (i. e. local blocks within the global dataset) is configured such that data is shared within node boundaries.
@@
Due to the plugin-based structure of PIConGPU, both the particle and radiation input required by MLapp are provided by distinct output plugins of PIConGPU, meaning that two parallel data streams are opened between PIConGPU and the MLapp. Each plugin prepares the current time step’s data and, upon having all data ready, ADIOS2 will gather the metadata, including remote read addresses to parallel rank 0 and provide the time step to the reader. The reader performs remote reads to its liking and closes the time step, indicating to the writer that the data can now be dropped.
@@
A unique and important requirement of the presented workflow is the ability to not only run but also scale on a large number of Frontier’s AMD MI250X GPUs. As our machine learning model is small enough to fit on a single GCD, parallel training of this model is done using data parallelism, where copies of the model are distributed across GCDs with each copy of the model receiving different chunks of data to train on. Once each model computes its gradients, all the instances of the model must do a collective all-reduce communication to average the gradients. We expect the scaling of the workflow to critically depend on the optimization of this all-to-all communication in PyTorch DDP, since communication within the PIConGPU simulation is only between next neighbors. We experienced a hard limit to scaling in our current setup: The all-to-all communication between PyTorch DDP ranks using the N/RCCL backend hits system limitations on the possible number of open sockets beyond 100 nodes. Potential options to circumvent this limit include using a libfabric backend for N/RCCL or using PyTorch DDP’s MPI backend. While the N/RCCL backend was able to support PyTorch up to the scale used in this study, we can already conclude that further extensive evaluation of the communication backends within PyTorch will become mandatory to provide a stable base for our near future scaling studies on Frontier. This area of research appears to be still under development, to the best of our knowledge [66].
@@ V-A Weak-Scaling
When training any kind of statistical model, in our case a deep neural network, one has to ensure that for each characteristic of the domain one hopes to learn, there are sufficient examples available in the dataset. When the dataset is provided in-transit, redundancy must still be built into the setup, either by using a large simulation volume or running multiple independent simulations concurrently, to allow the network to learn repeated patterns of the simulations and thereby distinguish relevant domain features, such as physical relationships, from noise. For this reason, in this work, we focus on a weak-scaling study of the proposed pipeline, as we anticipate this to be the most likely use-case. To evaluate parallel scaling as well as learning scaling, we performed simulations with in-transit training from 8 up to 96 nodes. The lower limit of 8 nodes is set by physical size constraints of the simulation. Fig. 8 shows the weak scaling behavior of the pipeline on Frontier. We define efficiency as the ratio of runtime at a specific size over runtime of the smallest size. Before averaging the times per training batch, we removed outliers more than four standard-deviations from the mean, as we observed single batches taking more than the mean time on Frontier.
While ideal scaling should correspond to 100 % efficiency across all sizes, communication overheads reduce efficiency in practice. Here, the efficiency of in-transit training reaches around 35 % at 96 nodes. We attribute the low parallel efficiency to two main reasons. The first reason is the inevitable all-to-all communication between PyTorch ranks that takes place to average the gradients during each backward pass and obtain a synchronized result on all ranks, which accounts for a deficit of , which is still a respectable result achieved by overlapping computations and communication in PyTorch DDP. The second reason is the computation of the two MMD loss terms, which amount to matrix dot products with data distributed across all ranks. As torch (less than version 2.2) did not offer a distributed primitive and the matrices involved are relatively small, even at the largest achieved scale, we opted for a naive implementation which replicates work between nodes. The global communication required here222torch.distributed.all_gather_into_tensor breaks the torch computational graph, i.e. synchronizes graph execution with host code at the invocation site. To our knowledge there is no high-level API available to stitch these operations into the graph to allow them to be executed asynchronously to both the host and each other, i.e no enqueueing of host functions or events.
@@ V-A1 Scaling Model Training
In our scaling runs, we explored training from scratch at large scale, sampling up to batches per sample consumed from the stream and found learning success up to about . We settled on a base-learning rate of , scaled according to the square-root scaling rule [60] depending on the actual number of GCDs. We observe a training speed up with larger batch sizes. That is, the increased learning rate overcompensates the training slow down due to larger batch sizes. Furthermore, it became evident that the three losses connected to the INN converged best at lower learning rates, while the VAE only found better minima at the highest learning rate, showing that separate learning rates and for the encoder/decoder block and the inversion block respectively need to be applied at large scales.
Within the scope of this study, we have to conclude that for in-transit training at very large batch sizes, hyper parameter studies do not transfer from small-scale experiments and thus have to be performed at scale. To enable in-transit learning at the scales we tested or beyond, comprehensive studies of the relations between the block learning rates and , batch sizes, and maybe even loss weights in Eq. (1) have to be performed. Exploring this parameter space for the large scale training requires the full study to be systematically performed at these scales (up to 384 GCDs), either by employing a streamed simulation or the same training data stored on disk. We argue that the latter is challenging at scale due to the lack of disk capacity and bandwidth. Efficient approaches to autotuning and optimization will be indispensable for this next step.
@@ V-B Quantifying the predictive capabilities of the machine learning method
The momentum prediction from the KHI shear surface was expected to be more difficult since particles from both directions cross in the shear surface, resulting in two separate and distinct particle populations. The ML reconstruction consistently predicts these two distinct populations. This is consistent with the interpretation that the INN has learned unsupervised that the vortex region undergoes more acceleration and thus emits more intense radiation than the bulk of the plasma. However, the momentum prediction is less accurate: the mean of (one of) these populations often does not match the PIConGPU data (Fig. 9(b vs. c) green). We suspect that this is partly due to our choice of region boundaries on the shear surface, which results in a much lower sampling rate of one flow direction compared to the other, leading to a lower accuracy in predicting the mean of the lower peak of the momentum distribution. It is also possible that this problem arises from the VAE’s decoder having difficulties reconstructing discontinuous distributions. Nevertheless, the successful prediction of shear regions indicates that the encoder extracts crucial features from the input data to recognize whether a region contains a vortex and encodes them in latent space.
The model clearly learned to partition the latent space into regions for different flow directions and vortex regions, which both the encoder and the inversion network learned to map to. These features are extracted using loss functions that do not impose physical rules and still encode physical properties of the system. Hence, in principle, they allow a simple, almost linear classifier to predict physical regimes. Evaluating such a classifier quantifies successful extraction of underlying physical rules.
We observed that main issue of our model is suboptimal reconstructions of vortex momenta, which we attribute to shortcomings of our decoder architecture, despite which the ML prediction still allows to visually identify KHI regions. Previous work on PC generation [67] observed that the latent space of an AE may contain more information than the decoder can reconstruct with satisfactory fidelity and that more sophisticated generative models, such as GANs or flows, can learn to produce more detailed reconstructions when conditioned on the same latent space.